# Intervening in Black Box: Concept Bottleneck Model for Enhancing Human Neural Network Mutual Understanding

### v1

@@ Appendix B Concept Replacement
The framework of the replacement method is illustrated in Figure X. It begins by identifying concepts that require intervention within each confusion class through concept intervention. These concepts are then ranked based on their frequency of occurrence, and the top $\bar{q}$ most frequently occurring concepts are selected for replacement. For each identified concept, positive concepts that do not require modification are also determined. Simultaneously, a replacement concept is selected from an external search set, with the objective of positioning it as far as possible from the embedding of negative concepts while bringing it closer to positive concepts. Cosine similarity is utilized to measure expression similarity, allowing us to score potential replacement concepts effectively.
Figure X: Concepts replacement framework. Intervention Concept Selection: Obtain the concepts that require intervention for each confused class. Intersection Sorting: The intervening concepts for all confused classes are intersected, and the concepts that need to be replaced are obtained by sorting the number of occurrences. Concept Replacement: Iteratively search for candidates in the External Concepts Search Set that are far away from the replacement concept but similar to other positive concepts of the same class, sort according to the $S_{replace}=S_{near}-S_{away}$, and select the optimal concept replacement.
Figure XI presents the accuracy comparison of NFResNet50 on Flower-102 before and after concept replacement. The results demonstrate that replacing different numbers of concepts leads to improved test accuracy for both the black-box baseline and the corresponding CBM.
Figure XI: Comparison of accuracy before and after the concept replacement of NFResNet50 and the corresponding CBM on Flower-102.
Appendix C Approximation and Intervention for CBM
We also conduct a class-based accuracy analysis on the approximate black-box CBM inference structure and compare it with the original black-box model. Figure XII presents a subset of accuracy comparison curves for confusing classes between the CBM-based approximate reasoning structure and the original black box. Additionally, Table I reports improvements achieved by CBM inference after applying a concept weight matrix in CBM-HNMU. The results indicate that the classification accuracy curve of CBM on intervention classes generally aligns well with that of the original black box. This suggests that CBM can effectively approximate the reasoning logic of the black box within locally confused classes. Furthermore, we evaluate the accuracy of the concept intervention algorithm based on CBM. The findings reveal that nearly all interventions help correct concept-related reasoning errors, leading to improvements in both class-level and global accuracy.
In Section 3.6, we use $\mathbf{pr}$ in Eq. (9) to transfer intervention knowledge while preserving the distribution of non-target classes. As a result, their accuracy remains largely unaffected. As suggested, we report the accuracy change on non-intervention classes for two models across three datasets in Figure XIII. The results show stable performance closely correlated with pre-intervention accuracy (scatter plot), along with a modest overall improvement (box plot).
@@ Appendix C Approximation and Intervention for CBM
Figure XII: Confusion class accuracy comparison curve between part of the CBM approximate reasoning structure and the original black box. The y-axis is the prediction accuracy of the class, and the x-axis is the corresponding intervention class ($\Gamma$).
Figure XIII: Before and after intervention in non-intervention classes.
Figure XIV: Visual related concepts masking. Intervention concepts and class name having the same color represent that the intervention concept belongs to the corresponding class, and the visual mask of samples belonging to this class is labeled according to these concepts.
Appendix D Detail for Intervention Concept Visual Masking
@@
These modified samples are then fed into the original black-box model to obtain new predictions and confidence scores, and compare with the original results. As shown in Figure XIV, in the first example, we visually masked $7$ misclassifies samples of sword lily on Flower-102 using NFResNet-50. Based on the vision-related concepts identified in the intervention’s natural language description, we masked the central pixels of the flowers in these samples.
@@
Appendix E Hardware and Software Settings
@@
Appendix F Parameter Settings

### v2

@@ Appendix A Concept Replacement
The framework of the replacement method is illustrated in Figure I. It begins by identifying concepts that require intervention within each confusion class through concept intervention. These concepts are then ranked based on their frequency of occurrence, and the top $\bar{q}$ most frequently occurring concepts are selected for replacement. For each identified concept, positive concepts that do not require modification are also determined. Simultaneously, a replacement concept is selected from an external search set, with the objective of positioning it as far as possible from the embedding of negative concepts while bringing it closer to positive concepts. Cosine similarity is utilized to measure expression similarity, allowing us to score potential replacement concepts effectively.
Figure I: Concepts replacement framework. Intervention Concept Selection: Obtain the concepts that require intervention for each confused class. Intersection Sorting: The intervening concepts for all confused classes are intersected, and the concepts that need to be replaced are obtained by sorting the number of occurrences. Concept Replacement: Iteratively search for candidates in the External Concepts Search Set that are far away from the replacement concept but similar to other positive concepts of the same class, sort according to the $S_{\mathbf{replace}}=S_{\mathbf{near}}-S_{\mathbf{away}}$, and select the optimal concept replacement.
Figure II presents the accuracy comparison of NFResNet50 on Flower-102 before and after concept replacement. The results demonstrate that replacing different numbers of concepts leads to improved test accuracy for both the black-box baseline and the corresponding CBM.
Figure II: Comparison of accuracy before and after the concept replacement of NFResNet50 and the corresponding CBM on Flower-102.
@@
Figure III: Confusion class accuracy comparison curve between part of the CBM approximate reasoning structure and the original black box. The y-axis is the prediction accuracy of the class, and the x-axis is the corresponding intervention class ($\Gamma$).
Figure IV: Before and after intervention in non-intervention classes.
Appendix B More Evaluation on Approximation and Intervention
We also conduct a class-based accuracy analysis on the approximate black-box CBM inference structure and compare it with the original black-box model. Figure III presents a subset of accuracy comparison curves for confusing classes between the CBM-based approximate reasoning structure and the original black box. Additionally, Table I reports improvements achieved by CBM inference after applying a concept weight matrix in CBM-HNMU. The results indicate that the classification accuracy curve of CBM on intervention classes generally aligns well with that of the original black box. This suggests that CBM can effectively approximate the reasoning logic of the black box within locally confused classes. Furthermore, we evaluate the accuracy of the concept intervention algorithm based on CBM. The findings reveal that nearly all interventions help correct concept-related reasoning errors, leading to improvements in both class-level and global accuracy.
To preserve the predicted distribution over non-target classes when transferring intervention knowledge, we apply the operator – $\mathbf{pr}$, to reassign only the residual probability mass from non-target classes to the intervention class, as defined in Eq. (9). Figure IV reports the accuracy shift for non-target classes across both models and all three datasets. The results show that performance is stable and closely aligned with pre-intervention accuracy (scatter plot), with a slight overall improvement (box plot).
Figure V: Visual related concepts masking. Intervention concepts and class name having the same color represent that the intervention concept belongs to the corresponding class, and the visual mask of samples belonging to this class is labeled according to these concepts.
Figure VI: More visualization of NFResNet50 on Flower-102.
Appendix C Intervention Concept Visual Masking
@@
These modified samples are then fed into the original black-box model to obtain new predictions and confidence scores, and compare with the original results. As shown in Figure V, in the first example, we visually masked $7$ misclassifies samples of sword lily on Flower-102 using NFResNet-50. Based on the vision-related concepts identified in the intervention’s natural language description, we masked the central pixels of the flowers in these samples.
@@
Appendix D Hardware and Software Settings
@@
Appendix E Parameter Settings
@@ Appendix G Limitations
Figure VII: More visualization of NFResNet50 on FGVC-Aircraft.
Figure VIII: More visualization of NFResNet50 on CUB-200.
Figure IX: More visualization of BotNet26 on Flower-102.
Appendix H Visualization and Explanation
We begin with additional visualizations of intervention-based explanations, as shown in Figures VIXIV. These illustrations highlight the relationship between natural language intervention concepts and changes in black-box visual representations before and after intervention. We present results using CBM-NHMU on NFResNet50, BotNet26, and RexNet100, with interventions applied to Flower-102, CUB-200, and FGVC-Aircraft. Notably, the samples are randomly selected from a subset where the black-box model’s original classification errors on the test set are corrected after intervention. Each corrected sample includes at least one pre- or post-correction class associated with the confused classes, ensuring a clear visual link to the intervention concept (see “Coverage" in the manuscript).
Figure X: More visualization of BotNet26 on FGVC-Aircraft.
Figure XI: More visualization of BotNet26 on CUB-200.
Figure XII: More visualization of RexNet100 on Flower-102.
In Figure VI, the baseline of NFResNet50 misclassifies sword lily as bougainvillea. The visually related intervention concept: "Prominent spadix in the center of the flower." prompts us to delete the related concept about the prominence in the center of flower to correct this error. We can find that the third most important visual concept extracted by the original black box before intervention corresponds exactly to this description, and the importance score of this concept is $0.11$. After the intervention, the black box gives a similar conceptual explanation, but the difference is that the importance score of the intervention concept dropped significantly ($\downarrow 0.06\rightarrow 0.05$). Combining the misclassifies sample image, we can also see that the visual features corresponding to the intervention concepts are indeed easy to confuse the two classes.
In Figure VII, NFResNet50 misclassifies BAE 146-300 as BAE 146-200. The visually related intervention concept: "Blue with a grey and white stripe." and "Large logo on the tail." prompts us to delete the related concept about the tail and the color-related feature of aircraft to correct this error. We can find that the second and the third most important visual concept extracted by the original black box before intervention corresponds exactly to the descriptions. After the intervention, we can find that the network relies less on the corresponding color-related feature and tail of the aircraft to classify real class. The concept importance scores of color-related features and aircraft tail dropped from $0.22$ to $0.18$ and $0.15$ to $0.13$, respectively.
Figure XIII: More visualization of RexNet100 on FGVC-Aircraft.
Figure XIV: More visualization of RexNet100 on CUB-200.
In Figure VIII, NFResNet50 misclassifies shiny cowbird as American crow. The visually related intervention concept: "Black webbed foot." prompts us to delete the related concept about the foot of bird to correct this error. We can find that the second and the third most important visual concept extracted by the original black box before intervention corresponds exactly to this description. After the intervention, we can find that the network relies less on bird footsteps to classify real class.
In Figure IX, BotNet26 misclassifies bougainvillea as siam tulip. The visually related intervention concept: "Very feminine shade of pink." prompts us to delete the related concept about the pink color of flower to correct this error. We can find that the first most important visual concept extracted by the original black box before intervention corresponds exactly to this description, and the importance score of this concept is $0.80$. After the intervention, the black box gives a similar conceptual explanation, but the difference is that the importance score of the intervention concept dropped significantly ($\downarrow 0.30\rightarrow 0.50$).
In Figure X, BotNet26 misclassifies Model B200 as DC-3. The visually related intervention concept: "Retractable tricycle landing gear." and "tricycle landing gear configuration with nose wheel steering." prompts us to delete the related concept about the gear of aircraft to correct this error. We can find that the second most important visual concept extracted by the original black box before intervention corresponds exactly to this description. After the intervention, we can find that the network relies less on gear of the aircraft to classify real class and the original score of the corresponding visual concept dropped from $0.24$ to $0.15$.
In Figure XI, BotNet26 misclassifies common tern as forsters tern. The visually related intervention concept: "A duck with a long, thin bill." prompts us to delete the related concept about the bill of bird to correct this error. We can find that the second and the third most important visual concept extracted by the original black box before intervention includes this description. After the intervention, we can find that the network relies less on bird bill to classify real class and the original score of the second most important visual concept dropped from $0.21$ to $0.09$. The third most important visual concept even disappears.
In Figure XII, RexNet100 misclassifies water lily as camellia. The visually related intervention concept: "heart-shaped spathe.", "curved spadix." and "prominent spadix in the center of the flower." prompts us to delete the related concept about the pistil of flower to correct this error. We can find that the second most important visual concept extracted by the original black box before intervention includes corresponding features. However, the stamen feature given in concept $2$ can easily be confused between the two classes. After the intervention, we clearly can find that the black box replaced concept $2$ with a more representative visual concept of the stamen to water lily.
In Figure XIII, RexNet100 misclassifies MD-87 as MD-80. The visually related intervention concept: "Large t-shaped tail fin.", "white with blue and red stripes.", "large tail fin with the airline’s logo.", etc. prompts us to delete the related concept about the tail with corresponding color-related and shape-related feature of aircraft to correct this error. We can find that the second most important visual concept extracted by the original black box before intervention corresponds exactly to this description. After the intervention, we can find that the network relies less on the tail of the aircraft to classify real class and the original score of the corresponding visual concept dropped from $0.45$ to $0.35$.
In Figure XIV, RexNet100 misclassifies white necked raven as American crow. The visually related intervention concept: "Head is grayish-brown." and "large, black bird with a long, thick beak." prompts us to delete the related concept about the head and beak of bird to correct this error. We can find that the first most important visual concept extracted by the original black box before intervention corresponds exactly to this description. After the intervention, we can find that the network relies less on bird head to classify real class and the original score of the corresponding visual concept dropped from $0.37$ to $0.28$.